OpenStudy (anonymous):

Challenge: Compute the are of the loop of curve \[y^2=x^2\frac{1+x}{1-x}\]

OpenStudy (anonymous):

Anyone

OpenStudy (anonymous):

Lets do brainstorming everyone give all ideas coming to mind

OpenStudy (anonymous):

integrate?

OpenStudy (anonymous):

I got it ans is \[2\int\limits_{-1}^{0}x^2\sqrt{\frac{1+x}{1-x}}\]

OpenStudy (anonymous):

\[2-\frac{\pi}{2}\]

OpenStudy (turingtest):

\[y=2\int_{-1}^{0}x\sqrt{\frac{1+x}{1-x}}dx\]isn't it? why is yours still x^2 squared?

OpenStudy (turingtest):

A=... sorry

OpenStudy (anonymous):

Sorry

OpenStudy (turingtest):

is that a no? is should be x^2

OpenStudy (turingtest):

or should not?

OpenStudy (anonymous):

not x^2

OpenStudy (anonymous):

turing yours was correct

OpenStudy (anonymous):

Should be

OpenStudy (turingtest):

\[A=\int_{-1}^{0}x\sqrt{\frac{1+x}{1-x}}dx\]\[u=1-x\]groovy :)

OpenStudy (turingtest):

but my sub wont' work

OpenStudy (anonymous):

that won't work.. i don't think

OpenStudy (anonymous):

Easier if you rationalise miltiply sqrt(1-x) in numerator and denominator

OpenStudy (turingtest):

oh yeah, ok

OpenStudy (anonymous):

I thought it was arc....I assume everyone else is looking at area...lol

OpenStudy (anonymous):

Sorry 1+x

OpenStudy (anonymous):

compute the area of the loop of the curve

OpenStudy (anonymous):

=area under curve

OpenStudy (anonymous):

Simplifies to \[\int\limits\limits_{-1}^{2}(\frac{x}{\sqrt{1-x^2}}-\sqrt{1-x^2}+\frac{1}{{\sqrt1-x^2}})dx\]

OpenStudy (turingtest):

where did the 2 come from? I have something else

OpenStudy (anonymous):

Typing error i menat it to be outside

OpenStudy (anonymous):

which should give the anwer approx.....using winplot .4292 But of course the mathematics is more fun than letting the computer do all of the work.

OpenStudy (turingtest):

ok, I still have a different expression though I'm trying to figure out if they're the same...

OpenStudy (turingtest):

am I doing something wrong?\[A=\int_{-1}^{0}x\sqrt{\frac{1+x}{1-x}}\cdot\sqrt{\frac{1+x}{1+x}}dx=\int_{-1}^{0}x\frac{1+x}{\sqrt{1-x}}dx\]\[=\int_{-1}^{0}\frac{x}{\sqrt{1-x}}+\frac{x^2}{\sqrt{1-x}}dx\]because my expression looks different ...or am I just tired?

OpenStudy (turingtest):

2 times all that*

OpenStudy (anonymous):

yeah (1-x)(1+x)=1-x^2 you look tired

OpenStudy (turingtest):

ok that's just a typo though

OpenStudy (turingtest):

I meant that the expression is different as far as I seeam I doing something wrong?\[A=2\int_{-1}^{0}x\sqrt{\frac{1+x}{1-x}}\cdot\sqrt{\frac{1+x}{1+x}}dx=2\int_{-1}^{0}x\frac{1+x}{\sqrt{1-x^2}}dx\]\[=2\int_{-1}^{0}\frac{x}{\sqrt{1-x^2}}+\frac{x^2}{\sqrt{1-x^2}}dx\]is what I meant

OpenStudy (anonymous):

If that is the case add 1and subtract 1 from x^2 to arrive at mine

OpenStudy (turingtest):

oh gotcha :P

OpenStudy (anonymous):

Nice!

OpenStudy (anonymous):

I feel double integrals no worth this problem it is not that difficult